Reports 1-1 of 1 Clear search Modify search
Infrastructures Monitoring (Infrastructure Monitoring)
fiori, tringali, paoletti, soldani - 10:57 Thursday 09 June 2022 (56097) Print this report
Assessment of maintenance works on AHUs

The following AHU maintenance works have been performed during the DAQ shut down week (11-15 April) and on May 17th:

  • AHU CleanRooms CEB (May 17th): replacement of belts and general internal and external filter and cleaning of the machine

  • AHU INJ Lab: checking and testing of motors, fans and their bearings, integrity of heat exchanger coils and cleaning of fins, internal and external cleaning of casing, replacement of steam cylinder and cleaning of humidifier pump and its filling tray, replacement of G3 pre-filters and F9 filters;

  • AHU DET Lab: check and verify motors, fans and their bearings, integrity of heat exchanger coils and cleaning of fins, internal and external cleaning of casing, replacement of steam cylinder and cleaning of humidifier pump and its filling tray, replacement of G3 pre-filters and F9 filters;

  • AHU Hall CEB: Checking and testing of motors, fans and their bearings, replacement of drive belts, integrity of heat exchanger coils and cleaning of fins, internal and external casing cleaning, replacement of G3 pre-filters;

  • AHU NEB and AHU WEB: Checking and testing of motor, fan and their bearings, replacement of drive belts, integrity of heat exchanger coils and cleaning of fins, internal and external casing cleaning, replacement of G3 pre-filters and F7 filters.

Comments to this report:
fiori - 12:34 Thursday 09 June 2022 (56098) Print this report

Noise assessment of maintenance works of CEB Clean Rooms AHU

Attached graphs compare the spectrum of Guralp CEB seismometer (vertical) before and after the maintenance works, respectively on midnight of May 17 and May 18.

Overall frequency lines associated to this AHU slightly moved in frequency (less than 0.1 Hz). Overall, we notice that some of the seismic lines associated to the belts frequency and harmonics have sliglity reduced in amplitude (up to a factor 2).

Images attached to this comment
Tringali - 14:26 Monday 20 June 2022 (56193) Print this report

At first glance, looking at Figure 1, the spectrum of the accelerometer, installed on the NEB AHU case, before (purple) and after (blue) maintenance is slightly changed. As shown in Figure 2, the seismic frequency lines are slightly shifted and the amplitude slightly increased.

The line at ~5.38 Hz should be due to the belts but this is different from the expected value (~4.19 Hz) and from that observed in the WEB AHU at ~4.29 Hz (the terminal AHU machines are identical), Figure 3.

Images attached to this comment
Paoletti - 15:29 Wednesday 06 July 2022 (56366) Print this report

Noise assessment of maintenance works of CEB Hall AHU

Attached is a comparison between the predictive accelerometers (ENV_CEB_AHU_RET_xxx) that were operational during maintenance works on the CEB HALL SUPPLY AHU system.

(Purple = before maintenance, Blue = after maintenance)

No major changes are visible, apart from a slight increase in frequency of some lines. The overall amplitude remain the same.

Images attached to this comment
Search Help
×

Warning

×